### v3

# Gen-C: Populating Virtual Worlds with Generative Crowds

Abstract
Over the past two decades, researchers have made significant steps in simulating agent-based human crowds, yet most efforts remain focused on low-level tasks such as collision avoidance, path following, and flocking. Realistic simulations, however, require modeling high-level behaviors that emerge from agents interacting with each other and with their environment over time. We introduce Generative Crowds (Gen-C), a generative framework that produces crowd scenarios capturing agent–agent and agent–environment interactions, shaping coherent high-level crowd plans. To avoid the labor-intensive process of collecting and annotating real crowd video data, we leverage large language models (LLMs) to bootstrap synthetic datasets of crowd scenarios. We propose a time-expanded graph representation, encoding actions, interactions, and spatial context. Gen-C employs a dual Variational Graph Autoencoder (VGAE) architecture that jointly learns connectivity patterns and node features conditioned on textual and structural signals, overcoming the limitations of direct LLM generation to enable scalable, environment-aware multi-agent crowd simulations. We demonstrate the effectiveness of Gen-C on scenarios with diverse behaviors such as, a University Campus and a Train Station, showing that it generates heterogeneous crowds, coherent interactions, and high-level decision-making patterns consistent with real-world crowd dynamics.
Keywords crowd simulation, multi-agent systems, data-driven method, large language models, variational graph auto-encoders
1 Introduction
Simulating human crowds has been an active area of research for several years, leading to a wide range of models and tools for generating realistic human behaviors [1, 2]. Several commercial solutions have also been integrated into game engines and rendering software to support large-scale crowd scenes [3, 4, 5, 6]. These approaches have advanced the realism of simulation, but they primarily target motion synthesis rather than the decision processes and interactions that govern behaviors over time.
At the algorithmic level, most methods have concentrated on low-level navigation tasks such as collision avoidance, path following, and group steering [7, 8]. While such methods effectively handle local interactions, they fall short of capturing the complex behaviors that arise when individuals pursue goals, interact socially, and respond to the affordances of their environment. For example, people stop to talk with friends, browse shop windows, or queue to buy tickets before boarding a train. Such behaviors need planning and coordination and are rarely modeled in current systems, which leads to repetitive dynamics that lack realism and adaptability.
In this work, we introduce Generative Crowds (Gen-C), a framework to automatically generate high-level behaviors of virtual agents, enabling multi-character simulations with diverse crowds. Throughout this paper, we use the term agent to denote a virtual character navigating within the crowd. At the core of our approach lies a crowd scenario graph: a time-expanded graph encoding both spatial and temporal evolution of the crowd, capturing high-level agent-to-agent and agent-to-environment interactions over time. With the advent of big data and deep learning, an obvious solution would be to learn such crowd scenario graphs from real data. Several data-driven methods have indeed modeled crowd by leveraging trajectories extracted from limited real-world datasets [9, 10, 11, 12, 13]. Nevertheless, the challenges and dependency on collecting and annotating data can quickly become a barrier and also limits the generalization capabilities of such solutions.
To break free from this dependency, we propose to automatically synthesize crowd datasets, featuring high-level behaviors, by leveraging the power of Large Language Models (LLMs) like the GPT framework [14]. While LLMs can in principle be used directly to script crowd scenarios, as we will show, this approach faces three main drawbacks: it scales poorly beyond a few agents, leading to repetitive behavior; it depends on heavy prompt engineering and domain expertise to produce coordinated behaviors; and it lacks efficiency for generating structured responses, requiring access to large and costly models. We therefore use LLMs only to create a small seed set of representative crowd scenarios that describe agent-agent and agent-environment interactions for given input scenes, which we subsequently represent as crowd scenario graphs.
Given such graphs, a generative model is trained to learn a distribution of crowd behaviors and interactions that we can then sample and synthesize new graphs. As most existing graph embedding algorithms are mainly developed for static graphs, they cannot capture the complex and dynamic nature of a crowd scenario graph. To that end, we propose a model architecture that leverages two Variational Graph Autoencoders (VGAEs) [15] to capture both how agents interact with each other and with their environment. The trained model can be conditioned on natural language, allowing users to populate scenes with structured crowds simulations directly from text.
Our approach builds upon recent advances in generative AI, where models have been successfully applied to text-driven human motion synthesis [16] and agent or robot planning [17]. However, these methods mainly focus on single-agent scenarios [18, 19, 20, 21, 22]. Closer to our work, text-based diffusion models have been explored for coordinating groups [23] and synthesizing crowd trajectories [24], but these methods do not capture structured, high-level behaviors. In contrast, Gen-C serves as a generative framework for high-level behavioral planning among multiple agents in a crowd, generating structured interaction graphs that model both agent–agent and agent–environment interactions, with a design extensible to text-guided generation. Overall, Gen-C introduces:
  1. 1.
    A graph-based representation that captures temporal agent-agent and agent-environment interactions.
  2. 2.
    A text-conditioned dual VGAE that jointly learns graph structure and features, enabling scalable generation of novel multi-agent scenarios.
  3. 3.
    A synthetic data pipeline, leveraging LLMs to bootstrap crowd scenarios, reducing reliance on real annotations.
2 Related Work
Crowd Simulation. Crowd simulation has a long history, with early works introducing rule-based models for navigation and flocking [25, 26]. Subsequent research proposed higher-level control mechanisms, such as crowd patches and sketch-based navigation fields [27, 28, 29], which allowed users to shape collective motion in structured environments. A major thrust of later work has been to increase realism by modeling heterogeneity, for example through group-level properties [30], individual personality traits [31, 32, 33], or stylistic variations across different types of crowds [34]. These methods highlighted the importance of diversity for believable simulations but typically required manual specification or extensive parameter tuning. In parallel, data-driven approaches emerged that learn behavioral patterns from real-world crowd trajectories [35, 36, 37], sometimes adopting graph-based representations [38, 39] that connect crowd modeling to broader developments in machine learning. More recently, reinforcement learning and hybrid imitation-learning methods [40, 13, 24, 12, 41] and diffusion-based generative approaches [42, 43] have demonstrated the potential of scalable and adaptive policies that capture richer dynamics. Despite these advances, most existing methods concentrate on low-level navigation or local diversity. They do not explicitly model structured, high-level multi-agent behaviors that depend on both agent-agent and agent-environment interactions over extended horizons. Our work addresses this gap by introducing a framework for generating heterogeneous, scenario-driven crowd behaviors that capture such multi-layered interactions.
Large Language Models and Autonomous Agents. LLMs have shown to be effective to domains beyond Natural Language Processing (NLP) such as task planning, embodied reasoning, autonomous agents, and floor plan design [44, 45, 46, 47, 48, 49]. Here, we explore an architecture that allows a LLM to generate a sequence of crowd behaviors, over time and space, that forms synthetic crowd datasets. We further exploit the power of generative NLP models to enable text-guided synthesis of high-level crowd behaviors. Along these lines a lot of work has explored generative models for text-driven human motion generation tasks [16]. Despite impressive results, including generating scene-and-language conditioned motions [18, 19, 20, 21, 22], such methods typically focus on animating individual characters. But a human crowd is much more than that, consisting of multiple agents performing a sequence of actions while interacting with other agents and the environment. Along these lines, the recent work of Ji et al. [23] uses a text-based diffusion model, in combination with LLMs, to steer crowds of agents but does not extend to other types of behaviors. Our work seeks to learn high-level behaviors captured via crowd scenario graphs that can complement the low-level steering capabilities of such agents.
Graph Generation. Beyond our domain, learning graph representations is a fundamental task across many real-world applications [50]. While much prior work has focused on learning static graphs using deep learning models such as VGAEs [15], recent research has extended these ideas to dynamic graphs, where the structure evolves over time. Approaches include neural temporal point processes and attention mechanisms for modeling time-dependent events [51], as well as deep architectures such as Temporal Graph Networks [52] and Spatio-Temporal Graph Convolutional Networks [53] for capturing temporal dependencies and predicting future states. While related, our work focuses on crowd scenario graphs: dynamic graphs with rich relational features that capture the evolution of crowds. Unlike most dynamic graph models aimed at tasks such as link prediction, node classification, or recommendation, our formulation targets the evolution of crowd interactions. Closely related, conditional graph generation has also gained attention, with most approaches generating only the graph structure while assuming fixed node and edge features. For instance, NGG [54] employs conditioned latent diffusion, GraphVAE [55] integrates conditional codes for guided generation, and MOOD [56] guides diffusion sampling with property gradients. Sequential models [57] use conditional inputs to initialize node and edge creation, while others embed conditional codes within message-passing schemes [58]. Complementary to these efforts, the Gen-C framework models both graph structures and node features via distinct, condition-aware latent distributions, enabling generation directly from textual input.
3 Overview
Preliminaries. A crowd scenario graph represents a crowd scenario, consisting of agents, a series of events, environment locations, and the associated actions and interactions. To simplify the generation and learning process, we define a finite set of actions and locations which, when combined, can describe a wide range of crowd behaviors. Specifically, an Action is a label from $\mathbf{\mathcal{A}ct}=$~{wait, sit, wander, queue, enter/exit, wave at, discuss, meet, object interact, talk on phone, read, look at, carry} that specifies an agent’s behavior at a given time. When multiple agents perform a shared action (e.g., discuss), this is referred to as an Interaction. An Environment Location is defined by its name, category, position, and scale. While the name may be abstract, the category is selected from a finite set: $\mathbf{\mathcal{L}oc}=$~{building, room, entrance, exhibit, furniture, outdoor area, item, service area}. Finally, an Event encodes a sequence of actions together with their location and participating agents. Specifically, each event consists of a set of action–location pairs (e.g., {discuss, room}), the participating agents, and a trigger condition that determines its initiation (i.e., starts after a specific event).
The Gen-C Framework. Our framework for generating crowd scenarios consists of three main components: Synthetic Data Generation, Learning, and Scenario Generation. First, we use a LLM to produce diverse synthetic scenarios through targeted queries that contains events describing agent actions, interactions, and environment placements. These scenarios are transformed into graph structures encoding action sequences, relationships, and interactions. Second, we train two synergistic VGAE models: one reconstructs the graph structure (agent interactions), while the other defines node features (agent actions and navigation). A text-conditioned prior integrates natural language, aligning the generated graphs with input descriptions. Finally, novel scenarios are synthesized by sampling from the learned latent distributions, producing contextually consistent multi-agent behaviors.
4 Synthetic Data Generation
Simulating diverse high-level crowd behaviors requires extensive high-quality data, yet collecting and annotating real crowds is costly and labor-intensive. To address this, we leverage recent advances in LLMs to synthesize data consistent with crowd dynamics.
4.1 Bootstrapping Crowd Scenarios via LLMs
To generate simulations that capture representative crowd dynamics, we design two tailored LLM queries, each serving a specific purpose. The process begins with a single-sentence input $S_{in}$ describing a crowd scenario, which serves as the seed for generation. We construct a set of $100$ such seeds and, for each, create $50$ paraphrased variants to increase diversity, yielding $5000$ textual inputs in total per theme; in this work we introduce two themes, University Campus and Train Station. Each sentence encodes three main components—a location phrase, one or more subject noun phrases, and one or more action verb phrases—without following a fixed syntactic template. We avoid rigid structures to promote linguistic diversity and reduce the risk od overfitting to narrow prompt patterns. An example of $S_{in}$ is “Around the campus park, students sit to have lunch, talk with others, and wave at passing friends.” Additional examples are provided in the supplementary material. We use OpenAI’s gpt-4.1 [59] and apply two sequential queries ($Q1$ & $Q2$), each constrained to return JSON-formatted outputs; exact prompts and parameters are listed in the supplementary material.
Generating the Environment ($Q1$). The first query prompts the LLM to generate a plausible environment layout from $S_{in}$, specifying Locations relevant to the scenario. For example, the output may include places like a “coffee shop”, “park”, or “entrance”, together with their category, position, and scale.
Authoring Crowd Events ($Q2$). The second query defines the sequence of agent actions and interactions, both with the environment and with other agents. It takes as input $S_{in}$ together with a simplified list of the previously generated environment, containing only the location names for context. The output is a list of events.
By combining responses from both queries, we construct a high-level crowd scenario $CS$ from the textual input $S_{in}$. We then iterate over the generated events to extract each agent’s actions and group memberships. As LLMs lack precise temporal detail, event sequences are derived from initiation constraints, and action durations are sampled from action-specific distributions. This yields a record $Rec_{i}$ for every agent $i$, capturing its actions and interactions throughout its lifespan in the simulated environment.
4.2 Graph-Based Representation
Training a generative crowd model requires an expressive data structure that captures agent–environment interactions and actions over time. We adopt a graph representation that meets these criteria and introduce the crowd scenario graph, a time-expanded graph modeling dynamic crowd scenarios through temporally encoded actions and interactions.
Figure 1: Snapshot of a crowd scenario graph. Nodes encode an action, a location, and an agent ID with timestep (i_ti\_t).
As described in Section 4.1, during simulation we collect $Rec_{i}$ for each agent $i$. Iterating through $Rec_{i}$, we create a node $V_{i}^{t}=(i,A_{i}^{t},L_{i}^{t})$, where $t$ is the current timestep, $A_{i}^{t}\in\mathrm{A}ct$, and $L_{i}^{t}\in\mathrm{L}oc$; the combination of $A_{i}^{t}$ and $L_{i}^{t}$ defines an agent-environment interaction at $t$. We set the action sequence of $i$ using a “sequence edge” $E_{i}^{t}=(V_{i}^{t},V_{i}^{t-1})$. Next, agent-agent interactions are captured by connecting nodes from different agents that share an action at a given timestep. A “share edge” $E_{i,j}^{t}=(V_{i}^{t},V_{j}^{t})$ is created for each pair of agents $i$ and $j$ interacting at $t$. At this stage, nodes of each agent $i$ are sequentially connected into a chain. If the chain has no “shared edges,” it forms an individual subgraph $H_{k}$ corresponding to agent $i$. When at least one “shared edge” exists, the chains of multiple agents merge into a subgraph representing a group of agents. Thus, for each $CS$, we construct an undirected graph $G=(V,E)$, consisting of multiple subgraphs $\mathcal{H}=\{H_{1},H_{2},\dots\}$, where $V$ is the set of nodes and $E$ the set of edges. Finally, each subgraph $H_{k}\in\mathcal{H}$ is treated as an individual training sample $(S_{in},H_{k},\text{agents}_{k})$, where $H_{k}$ is the subgraph for group $k$, and $\text{agents}_{k}$ the agents in that group. Figure 1 presents a snapshot of a subgraph with five agents from a crowd scenario graph.
5 Learning Crowd Scenario Graphs
While a synthetic LLM-generated crowd scenario graph, can drive a crowd of simulated agents, our goal is to generalize beyond manually querying an LLM whenever input conditions change or the simulation scale increases. To achieve this, we train a generative model on a dataset of crowd scenario graphs, enabling the synthesis of novel graphs directly from learned distributions.
5.1 Data Preprocessing
For each sample $(S_{in},H_{k},\text{agents}_{k})$, we treat each subgraph $H_{k}$ as an individual undirected graph $G_{k}=(V_{k},E_{k})$, where $n=|V_{k}|$ is the number of nodes and $m=|E_{k}|$ the number of edges. The adjacency matrix $\mathcal{A}_{G_{k}}\in\mathbb{R}^{n\times n}$ is symmetric and encodes edge information: entries are 0 if there is no edge between $v_{i}$ and $v_{j}$, $1$ if the edge is of type “sequence,” and 1-1 if it is a “share” edge.
Node Ordering. We enforce a canonical ordering of nodes to ensure consistency when constructing adjacency matrices. Different permutations of node indices can lead to inconsistency and higher training complexity, particularly for graph reconstruction tasks. To address this, we define a custom ordering where nodes are sorted first by event index, then by agent identifier (ID), and finally by the sequence index within the agent’s actions, ensuring a consistent representation across samples. Fig. 2 shows sample adjacency matrices with canonical and arbitrary orderings.
Figure 2: Canonical ordering (top) vs. arbitrary ordering (bottom). We sort nodes by event, agent ID, and timestep.
Node Features. Each node is annotated with a dense feature vector comprising the action, location, agent ID, normalized sequence index within the agent’s timeline, global sequence index, and Laplacian [60] positional encodings. Categorical attributes (action, location, agent ID) are mapped to learnable embeddings of dimensions $16$, 88, and $16$, respectively. The normalized time index adds a scalar feature, global index uses a 4D sinusoidal embedding, and the top-4 eigenvectors of the normalized Laplacian add structural context. Concatenating all components yields a $49$-dim node representation.
Graph-Level Conditioning. Beyond node-level features, we compute global signals that capture the context of a crowd scenario and serve as conditioning inputs for our model (see Section 5.2). Specifically, we include the normalized counts of nodes, agents, and events in each graph, using maximum values of $100$, $15$, and $30$ respectively. We also calculate an action-frequency vector that summarizes the distribution of action classes. These scalars and the action-frequency profile are concatenated with the embedding of $S_{in}$, obtained from the pretrained SentenceTransformer all-mpnet-base-v2 [61], and jointly form the condition vector $C$.
Figure 3: We use two synergetic VGAEs—one for graph reconstruction and one for node feature reconstruction—whose encoders map inputs to latent spaces regularized by a shared Condition Network.
5.2 Training
We train two conditional VGAEs that jointly learn to reconstruct graph structure and node features from crowd scenario graphs; we present the model’s architecture in Fig. 3. The two models share a common encoder that produces latent embeddings, while using distinct decoders specialized for structure (VGAE-S) and features (VGAE-F). The encoder learns an approximate posterior distribution, $q(Z\mid X,C)$, where $Z_{S}$ and $Z_{F}$ are graph-level latents for structure and features, respectively. We observe that using a fixed standard normal prior, $p(Z)=\mathcal{N}(0,I)$, leads to a mismatch between $q(Z\mid X,C)$ and $p(Z)$, resulting in posterior collapse that degrades generalization. To mitigate this mismatch, we introduce a prior network that parameterizes the conditional priors $p(Z_{S}\mid C)$ and $p(Z_{F}\mid C)$, where $C$ is the condition vector. This conditional prior aligns the posterior with contextual and structural cues, enabling generation of graphs consistent with given conditions.
Encoder. Both VGAE-S and VGAE-F share the same encoder architecture, which maps an input graph $G=(V,E)$ with adjacency $\mathcal{A}_{G}$ and node features $\mathbf{X}_{G}$ into a latent representation. The encoder stacks multiple GINE layers [62], each implemented as a message-passing step followed by a residual connection. At layer $k$, node embeddings are updated as:
$\mathbf{h}_{v}^{(k)}=\mathbf{h}_{v}^{(k-1)}+\text{GINE}^{(k)}\!\big(\mathbf{h}_{v}^{(k-1)},\{(\mathbf{h}_{u}^{(k-1)},e_{uv}):u\in\mathcal{N}(v)\}\big),$ (1)
where $\mathbf{h}_{v}^{(k)}$ is the representation of node $v$, $\mathcal{N}(v)$ its neighbors, and $e_{uv}$ the edge feature (“sequence”/“share”). After message passing, node embeddings are aggregated with global add pooling into a graph-level vector, normalized, and concatenated with the condition embedding $C$. Fully connected layers project this joint representation into the mean and log-variance of the Gaussian posterior,
$q(Z\mid X,C)=\mathcal{N}\!\big(\mu(X,C),\operatorname{diag}(\exp(\log\sigma^{2}(X,C)))\big),$ (2)
where $\log\sigma^{2}(X,C)$ is the predicted log-variance and $\exp(\log\sigma^{2})$ recovers a diagonal covariance with entries given by per-dimension variances. This parameterizes the latent distributions $q(Z_{S}\mid X,C)$ and $q(Z_{F}\mid X,C)$ for structure and features, respectively.
Structure Decoder. The structure decoder reconstructs the adjacency matrix $\widehat{\mathcal{A}}_{G}$ from the latent variable $Z_{S}$ and condition vector $C$. The concatenated vector $Z_{S};C]$ is processed by an MLP with nonlinear activations, and its outputs are mapped through a $\tanh$ function to produce edge values in [1,1][-1,1]. These values are assembled into a symmetric adjacency matrix,
$\widehat{\mathcal{A}}_{G}=U+U^{\top},\quad U_{ij}=\tanh\!\big(f_{\theta}([Z_{S};C])_{ij}\big)\;\;\text{for }i<j,$ (3)
with a zero diagonal. The sign of each entry encodes the edge type (“sequence” or “share”), while the magnitude reflects confidence.
Feature Decoder. The feature decoder reconstructs node features $\mathbf{X}_{G^{\prime}}$ from the latent variable $Z_{F}$, the condition vector $C$, and graph-level positional encodings. At the node level, $Z_{F}$ and $C$ are expanded and concatenated with projected positional encodings, which combine Laplacian embeddings from the share-only subgraph (capturing the connectivity in individual groups induced by “share” edges) and the full graph (capturing the overall structure), together with a normalized group identifier. The fused representation is processed by stacked GINE layers with pre-normalization and residual connections, enabling message passing over the graph.
To enforce consistency among agents that interact, node embeddings are collapsed into connected components defined by share edges, ensuring that members of the same group, at same timestep, are assigned a common action and location. Class logits are predicted at this group level and then uncollapsed back to the node level for supervision. Formally, for each node $v\in V$, the decoder outputs categorical distributions over actions and locations,
$\widehat{\mathbf{X}}^{\text{act}}_{v}=\text{softmax}(f^{\text{act}}_{\theta}(h_{v})),\qquad\widehat{\mathbf{X}}^{\text{loc}}_{v}=\text{softmax}(f^{\text{loc}}_{\theta}(h_{v})),$ (4)
where $h_{v}$ is the reconstructed embedding of node $v$ after the collapse–predict–uncollapse pipeline. These distributions are directly compared with the ground-truth labels $\mathbf{X}_{v}^{\text{act}}$ and $\mathbf{X}_{v}^{\text{loc}}$ during training.
Condition Network. Both VGAE models use the same condition network, which encodes the text embedding of $S_{in}$ together with normalized counts of nodes, agents, and events, as well as an action-frequency vector, to produce the condition vector $C$. This vector is also used to train two prior networks that parameterize the conditional priors $p(Z\mid C)$ for VGAE-S and VGAE-F, yielding their respective means $\mu_{\text{prior}}$ and log-variances $\log\sigma_{\text{prior}}^{2}$.
The condition vector is integrated into both the encoding and decoding phases. In the encoder, $C$ is concatenated with the pooled graph representation to produce the posterior distribution $q(Z\mid X,C)$, parameterized by $\mu$ and $\log\sigma^{2}$. During decoding, it is concatenated with the latent variable to guide reconstruction.
Training Objective. The two VGAE models are trained by maximizing the evidence lower bound (ELBO) [63], which balances reconstruction fidelity with regularization of the latent space. For the graph-structure model, VGAE-S, the reconstruction loss $\mathcal{L}_{S}$ compares the predicted adjacency $\widehat{\mathcal{A}}$ against the ground-truth adjacency $\mathcal{A}$ using a smooth L1-error loss (SmoothL1):
$\mathcal{L}_{S}=\sum_{i,j}\text{SmoothL1}\!\left(\mathcal{A}_{ij},\widehat{\mathcal{A}}_{ij}\right).$ (5)
For the feature model (VGAE-F), the reconstruction loss $\mathcal{L}_{F}$ compares categorical predictions of actions and locations with ground-truth labels using cross-entropy (CELoss) at the node level:
$\mathcal{L}_{F}=\sum_{v\in V}\Big(\gamma\text{CELoss}\big(\mathbf{X}_{v}^{\text{act}},\widehat{\mathbf{X}}_{v}^{\text{act}}\big)+\text{CELoss}\big(\mathbf{X}_{v}^{\text{loc}},\widehat{\mathbf{X}}_{v}^{\text{loc}}\big)\Big),$ (6)
where $\gamma=1.5$ is a balancing term. In both models, the latent variables are regularized with a KL divergence term that penalizes deviation of the posterior from the conditional prior:
$\mathcal{L}_{KL}=\mathbb{E}_{z\sim q(Z\mid X,C)}\left[\log q(Z\mid X,C)-\log p(Z\mid C)\right].$ (7)
The final training objectives are:
$\mathcal{L}_{VGAE-S}=\mathcal{L}_{S}+\beta_{S}\,\mathcal{L}_{KL_{S}},\qquad\mathcal{L}_{VGAE-F}=\mathcal{L}_{F}+\beta_{F}\,\mathcal{L}_{KL_{F}},$ (8)
where $\beta_{S}$ and $\beta_{F}$ are annealing factors that control the relative weight of the KL divergence. Both are scheduled using a cyclical annealing strategy [64], in which the KL weight is periodically ramped from zero to its maximum value over the course of training.
6 Crowd Scenario Generation
At inference time, novel crowd scenarios are generated by sampling from the learned priors using the input condition $C$. The condition includes: a textual description $S_{in}$, graph-level statistics (counts of nodes, agents, and events), and a frequency map of the available actions. From this condition, the structure latent $Z_{S}\sim p(Z_{S}\mid C)$ and the feature latent $Z_{F}\sim p(Z_{F}\mid C)$ are drawn and decoded into an adjacency matrix $\widehat{\mathcal{A}}$ and node features $\widehat{\mathbf{X}}$, respectively. This ensures that both connectivity and node attributes are contextually aligned with the full conditioning signal.
Each generated graph corresponds to a subgraph that encodes the actions and interactions of a group of agents, as introduced in Section 4.2; the framework scales naturally to arbitrary numbers of agents or groups. From the decoded graphs, we extract per-agent sequences of actions, locations, and interactions. To instantiate these scenarios in a virtual scene, the environment’s layout can either be provided manually or synthesized with a lightweight LLM query ($Q1$) (see Section 4.1). The resulting scenarios are simulated and visualized in Unity [5], with examples shown in Fig. 8 and additional animations included in the supplementary video.
7 Experiments and Evaluation
We train our model on two synthetic datasets capturing crowd scenarios in a University Campus and a Train Station. Using the bootstrapping procedure described in Section 4.1, we generate 500 scenarios per dataset. These are converted into interaction graphs as outlined in Section 4.2, yielding 22.922.9k subgraphs for University Campus and 25.125.1k for Train Station. We split each dataset into 80% training, 15% validation, and 5% testing. The code, datasets and trained models will be released upon acceptance.
For model training (cf. Fig. 3), the encoder consists of two GINE layers with hidden dimension 128, each followed by dropout ($p=0.2$). Both VGAE-S and VGAE-F use a latent dimension of 32. The graph-structure decoder is a three-layer MLP with hidden dimension 128, while the feature decoder applies two GINE layers (hidden dimension 128) followed by two linear heads that map to 13 and 8 output classes for actions and locations, respectively. The condition network projects text embeddings (768D), scalar values (3D), and action frequencies (13D) into 128 dimensions each; their concatenation is passed through an MLP to produce a 32-dimensional condition vector; the prior network maps this to the parameters of the prior distribution via two linear layers. LeakyReLU activations are used throughout, with LayerNorm applied in the encoder, both decoders, and condition network for stability. Models are trained for 500 epochs with a batch size of 128, learning rate of $5\times 10^{-4}$, linearly decayed to $1.25\times 10^{-4}$ over training, and Adam optimizer with weight decay of $3\times 10^{-4}$. The $\beta$ coefficient is annealed cyclically over 100-epoch cycles, in the ranges $[0.001,3]$ for VGAE-S and $[0.001,1]$ for VGAE-F. Training each model required approximately one hour on a single NVIDIA RTX 4070 Ti GPU.
7.1 Ablation Study
To evaluate the impact of our design choices, we compare the proposed Gen-C model against two variations of our architecture on the test set: (i) a version trained without canonical node ordering (w/o Can.), and (ii) a single VGAE model jointly trained on both structure and features (Single). Additionally, we include a constrained random baseline (Rand.), designed to support fair comparisons by generating graphs that conform to the same structural constraints as the ground truth (e.g., each node has 1–2 temporal edges). Node actions and locations are sampled from empirical priors and perturbed with mild noise, preserving marginal distributions while introducing controlled variability.
Performance is reported using the KL divergence (KLD) between the distributions of ground-truth and generated samples across both structural and semantic measures. Lower values indicate closer alignment between generated and real samples. Structural fidelity is assessed using standard graph descriptors [65]: the degree of each node $v$ is its number of incident edges; the clustering coefficient (CC) quantifies local transitivity, defined as $c(v)=\tfrac{2T(v)}{k_{v}(k_{v}-1)}$ where $T(v)$ is the number of triangles involving $v$ and $k_{v}$ its degree (with $c(v)=0$ if $k_{v}<2$); the graph diameter is the maximum shortest-path length between connected nodes; and the average path length (APL) is the mean shortest-path distance across all connected node pairs. Semantic alignment is evaluated by comparing the distributions of predicted and ground-truth action and location labels.
As shown in Table 1, Gen-C consistently achieves the lowest divergence across most metrics in both datasets. The variant without canonical ordering performs poorly on structural measures, highlighting the importance of consistent node indexing for stable learning. The single VGAE model improves over this baseline but underperforms Gen-C, especially on semantic metrics, suggesting that separating structure and feature learning yields more accurate alignment between agent interactions and their actions/locations; interestingly, the diameter statistic benefits from joint training. The random baseline, while constrained, performs poorly on graph-structure metrics but shows respectable feature-level alignment, since actions and locations are sampled from the provided input distributions. Overall, the ablation confirms that the full Gen-C framework is necessary to faithfully capture both the structural and behavioral properties of crowd scenarios.
Table 1: KL divergence (KLD) to ground-truth distributions.
Metrics University Campus (KLD) $\downarrow$ Train Station (KLD) $\downarrow$
w/o Can. Single Rand. Gen-C w/o Can. Single Rand. Gen-C
Degree 3.496 0.062 0.595 0.056 3.289 0.051 0.468 0.033
$C$ 0.955 0.034 0.221 0.031 0.147 0.053 0.438 0.038
Diameter 3.463 0.145 0.494 0.156 3.649 0.094 0.804 0.104
APL 0.923 0.276 0.885 0.147 0.714 0.144 1.235 0.099
Action 0.180 0.358 0.225 0.177 0.128 0.340 0.368 0.106
Location 0.024 0.058 0.024 0.017 0.019 0.031 0.047 0.018
7.2 Assessment of LLM-bootstrapped Data
Gen-C leverages LLMs to synthesize crowd scenario graphs, thereby reducing the dependence on real-world labeled data. To evaluate this approach, we annotated high-level action plans of individuals interacting in Train Stations, using a curated set of YouTube videos totaling approximately 30 minutes of footage. Comparable data for university campus environments are scarce, primarily due to ethical restrictions on video collection, and existing datasets such as UCY Dataset [36] exhibit limited diversity in observed actions. Specifically, we divide each video clip into $5$-second windows and assign an action label from $\mathbf{\mathcal{A}ct}$ to each person visible in the frame.
[…]
Figure 4: Per-agent action transitions (LLM) and deviation from Real ($\Delta$) in Train Station; rows“from”, columns “to”.
As an additional statistic, we measure diversity of action plans using the normalized entropy of per-agent action sequences. Let $\mathcal{S}$ be the set of distinct action sequences and $p(s)$ their empirical frequencies. We compute $H(p)\;=\;-\sum_{s\in\mathcal{S}}p(s)\,\log p(s)$ and report the normalized value $\hat{H}=H(p)/\log|\mathcal{S}|\in[0,1]$: $\hat{H}=0$ means one pattern dominates; $\hat{H}=1$ means all patterns are equally likely. By this measure, LLM sequences closely match the annotated real-world data: $\hat{H}_{\text{LLM}}=0.949$ $(95\%\ \text{CI }[0.916,\,0.970])$ vs. $\hat{H}_{\text{Real}}=0.936$ $(95\%\ \text{CI }[0.935,\,0.939])$. The overlapping intervals indicate that LLM data follow real-world sequence variety for the given context.
7.3 Rationale for a Dedicated Learned Generator
Scalability, Validity, and Cost. We compare vanilla LLM generation against our learned generator (Gen-C) on University Campus under increasing agent counts $A_{count}\in[20,160]$. For consistency we proportionally scale the requested event range and the LLM’s max response tokens, and evaluate $50$ seeds per setting while reusing the same $S_{in}$ and a fixed environment designed per seed. The LLM uses the $Q2$ query (Section 4.1), updating agents/events/tokens per setting; Gen-C samples subgraphs from its prior and composes them until the target $A_{count}$ is reached. For each run we report four metrics: (i) action-sequence entropy $H_{\text{seq}}$ (normalized entropy over per-agent action sequences), (ii) inference time (LLM: end-to-end completion; Gen-C: sampling + validation over composed subgraphs), (iii) token response usage (LLM only), and (iv) validity (LLM only) as the drop rate, i.e., the fraction of seeds whose realized agent count falls outside a tolerance $[0.8,1.2]$ of the target. In Fig. 5, we report means for metrics with $95\%$ confidence intervals.
As $A_{count}$ grows, the LLM shows a monotonic decrease in $H_{\text{seq}}$ (agents follow increasingly similar plans), a rising drop rate (more failures to meet the target $A_{count}$, and sharply increasing token usage and latency. In contrast, Gen-C maintains stable action sequence diversity and low, slowly increasing inference time. We also tested smaller LLMs (gpt-4.1-mini and gpt-4.1-nano), which exhibited high drop rates ($0.32$ and $0.94$, respectively) even at baseline $A_{count}=20$. Overall, we show that a dedicated learned generator preserves diversity and reliability at scale, whereas prompt-based LLM generation degrades in both quality and efficiency under the same conditioning, revealing the need of a learned generator.
Figure 5: vanilla LLM vs Gen-C under increasing agent scale.
Prompt Engineering. LLM-based crowd scenario generation currently relies on hand-crafted rules and constraints, creating a heavy configuration burden for non-experts. We employ a schema-guided query enforcing hard constraints (e.g., agent/event limits, diversity, no overlaps). Achieving validity requires (i) detailed instructions with few-shot examples, (ii) a custom JSON schema with post-validators, (iii) tuning of decoding parameters, and (iv) stepwise prompting with retries. These settings are brittle: small changes in scene scale and context often demand re-tuning, making the process error-prone; prompt details are in the supplementary material. In contrast, our learned generator needs only a short description and a few intuitive scalars (#agents, #nodes, #events, action-frequency priors). Without prompt engineering, it sustains high validity, fidelity, and coverage with predictable latency and cost—enabling novice users to easily populate virtual scenes with diverse crowds.
7.4 Quantitative Results
Reconstruction Quality. We test the reconstruction quality of Gen-C by sampling from the conditional prior $p(Z\mid C)$ on testing set, using the input conditions of each ground-truth scenario. Our evaluation considers both (i) graph structure and (ii) node features. For structure, we report standard graph statistics, as previously introduced in Section 7.1. For features, we compare the distributions of action and location labels from $\mathcal{A}ct$ and $\mathcal{L}oc$.
Figure 6: Graph Structure and Feature reconstruction quality from sampling using the conditional prior on unseen data.
As shown in Fig. 6, we compute the KLD between ground-truth and reconstructed graph distributions; lower values indicate better alignment. VGAE-S accurately recovers structural statistics, while VGAE-F achieves strong feature reconstruction with label distributions close to the ground truth. Performance is consistent across datasets with different dynamics. For example, in the Train Station dataset “wait” is more common and “meet” less so compared to University Campus, reflecting context-specific behaviors such as travelers waiting for trains versus students meeting on campus. Similarly, “service station” location appears more frequently in Train Station, corresponding to information booths or arrival/departure boards, whereas outdoor areas are more prevalent in the University Campus scenario. Additionally, “object interact” and “discuss” stand out in both datasets; the former capturing context-specific actions (e.g., eating in a campus cafeteria or grabbing a ticket at a station), and the latter reflecting a common social behavior.
Latent Space Analysis. We assess how closely graphs sampled from the model prior match the training distribution using a fixed evaluator: encoders trained only on training data. For each theme (University Campus & Train Station), we embed the training set $T$ and predicted set $P$ into latent spaces $Z_{T}$ and $Z_{P}$, derived from encoder posterior means $\mu$ for structure and features. We compute the Fréchet Inception Distance (FID) [66] and Maximum Mean Discrepancy (MMD) [67]: FID compares first- and second-order moments of the latent distributions, while MMD captures higher-order and local differences via a kernel-based similarity. Together, they provide complementary measures of real-generated space alignment.
$\mathrm{FID}$ $=\lVert\mu_{T}-\mu_{P}\rVert_{2}^{2}+\mathrm{Tr}\!\left(\Sigma_{T}+\Sigma_{P}-2(\Sigma_{T}\Sigma_{P})^{1/2}\right),$ (9)
$\mathrm{MMD}^{2}$ $=\mathbb{E}_{TT}[k_{\gamma}(x,x^{\prime})]+\mathbb{E}_{PP}[k_{\gamma}(y,y^{\prime})]-2\,\mathbb{E}_{TP}[k_{\gamma}(x,y)],$
where $k_{\gamma}(x,y)=\exp(-\gamma\|x-y\|_{2}^{2})$.
Table 2: Evaluator-space metrics for real vs. sampled graphs.
[…]
Uni. Campus Structure 1.950 0.078 2.921 0.099
Features 9.286 0.032 17.578 0.091
Train Station Structure 0.534 0.058 3.979 0.076
Features 7.674 0.030 16.004 0.109
As shown in Table 2, FID and MMD remain low (FID: 0.53–9.29; MMD: 0.030–0.078) across both themes, with Train Station aligning closer to the training distribution than University Campus. The structure model yields lower FID and feature model lower MMD: structure latents match global moments, while feature latents preserve local neighborhoods. We also run cross-domain checks by mixing model/data: the model trained on University Campus generates samples on Train Station and vice versa. As reported in Table 2 (Mix), both FID and MMD degrade, reflecting reduced alignment across domains; this confirms that Gen-C learns appropriate latent spaces, capturing unique structure and dynamics of each domain.
Figure 7: 2K True/Predicted latent vectors for Train Station.
Moreover, in Fig. 7, we plot 2K true and predicted latent vectors via PCA for VGAE-S (Structure) and VGAE-F (Features) on Train Station. In-domain, true and predicted latents largely co-locate. We also show cross-domain “mix-and-match” (Mix) projections, which exhibit reduced overlap and clear domain shifts, consistent with degraded FID/MMD in Table 2. The mixed-structure space retains some overlap, as similar structures span themes, whereas the feature latent space shows a strong mismatch, indicating that node features are tightly coupled to input context. Overall, both metrics and embeddings suggest our generator produces crowd scenario graphs well aligned with the training distribution in evaluator space.
7.5 Qualitative Results
In Fig. 8, we show rendered results across both datasets. Gen-C reproduces sequences of behaviors observed in real environments, including agents queuing for food or coffee, reading announcements, waiting for trains, chatting over lunch, and wandering through open areas. These examples highlight the model’s ability to generate heterogeneous yet coherent multi-agent actions and interactions. Additional qualitative results and longer sequences are provided in the supplementary video.
Figure 8: Generated render results from our framework.
8 Discussion and Future Work
We introduce Gen-C, a framework for populating virtual scenes with multiple agents that perform diverse, high-level actions and interactions from a short textual prompt. Gen-C couples a schema-guided LLM with two learned latent spaces (graph structure & features) to synthesize crowd scenario graphs that drive agents behavior. In contrast to traditional crowd systems focused on local navigation (e.g., goal seeking and collision avoidance), Gen-C targets the semantic layer—modeling activities, locations, and social interactions—to generate plausible, varied, and scalable scenarios.
Our current implementation does not support modeling long-term agent intentions, agents cannot switch actions mid-execution, and action durations are sampled from manually defined distributions. Additionally, behaviors are constrained to a predefined list of high-level actions, requiring retraining for new actions.
For future work, we aim to equip agents with a memory module, ensuring a more natural progression of actions. We would also like to combine Gen-C with a crowd simulator, allowing high-level behaviors to control low-level navigation policies, creating more dynamic and plausible simulations. Furthermore, we plan to explore combining latent spaces from different datasets, such as indoor/outdoor scenarios, using techniques such as distillation to enable lifelong learning and dynamic simulation adaptability. We envision that these will allow us to populate virtual worlds with diverse human-like characters, paving the road towards automatic population of virtual worlds.